# A Grey-box Attack against Latent Diffusion Model-based Image Editing by Posterior Collapse

### v3

@@ Introduction
Objective 1: $\min_{\bm{\delta}}\mathcal{D}(f(\bm{x}+\bm{\delta}),\bm{x}+\bm{w})\,\,\,s.t.\,\|\bm{\delta}\|_{p}\leq\epsilon,$
Objective 2: $\max_{\bm{\delta}}\mathcal{D}(f(\bm{x}+\bm{\delta}),f(\bm{x}))\quad s.t.\,\|\bm{\delta}\|_{p}\leq\epsilon,$
where $f(\cdot)$ is a kind of LDM-based image editing method; $\bm{w}$ refers to a kind of watermark artifact; $\mathcal{D}(\cdot)$ measures the perceptual distance between two inputs, indicating the visual consistency of two images in human visual perspective; $\|\cdot\|_{p}$ means applying a constraint on the vector, in most cases, this serves to maintain the visual integrity of the adversarial sample, following the $\ell_{\infty}$ norm. Notably, Objective 1 will cause similar results to Objective 2, i.e., there will be some distance in $\mathcal{D}$’s space between $f(\bm{x}+\bm{\delta})$ and $\bm{x}$.
Existing methods in the literature typically address either Objective 1 or Objective 2. However, both approaches often require extensive white-box information about the target model, particularly access to the neural backbone U-Net of the LDM. This heavy reliance on model-specific details limits their transferability and applicability across different LDM architectures, requiring more computing resource.
Our method focuses primarily on Objective 2, but takes a fundamentally different approach. Instead of relying on detailed knowledge of the entire LDM pipeline, we exploit the inherent characteristics of LDM-based editing by targeting the VAE component, which is common across various LDM architectures. This strategy allows us to achieve the goal of maximizing the disparity between $f(\bm{x}_{adv})$ and $f(\bm{x})$ without requiring extensive access to model-specific information, particularly the compute-intensive and model-specific U-Net component. By concentrating on the VAE, our approach aligns more closely with real-world scenarios where full model access may not be available, providing an efficient solution to prevent infringers from exploiting LDM-based image editing outputs.
Posterior Collapse
Variational Autoencoder.
A key observation driving our approach is the ubiquity of VAEs in the architecture of LDMs. VAEs serve as a foundational component across different LDM implementations, often with only minor variations between models. For instance, the VAE used in Stable Diffusion 2.0 is essentially a fine-tuned version of the one employed in Stable Diffusion 1.5. This commonality presents a strategic opportunity: by focusing on the VAE, we can potentially affect a wide range of LDMs without requiring detailed knowledge of their specific architectures.
@@
By general consensus, training of VAE usually suffers an optimization issue called posterior collapsing (Razavi et al. 2019), that is the KL divergence term of Equation 5 dominates the overall loss such that all posteriors collapse to the uninformative prior and the reconstruction error or the likelihood term is ignored. We draw inspiration from this vulnerability and aim to propose adversarial attacks to cause the posterior collapse during inference stage.
Posterior Collapse Loss Function.
@@
We aim to leverage the divergence measure between $q(\bm{z}|\bm{x})$ and a target distribution $p^{\ast}(\bm{z})$. The objective is to generate $\bm{x}_{adv}$ by minimising the KL divergence between $q(\bm{z}|\bm{x})$ and $p^{\ast}(\bm{z})$. The KL divergence of two multivariate Gaussian distributions is:
$\begin{split}D_{\text{KL}}({\mathcal{N}}_{1}\parallel{\mathcal{N}}_{2})&=\frac{1}{2}[\ln\frac{|\bm{\Sigma}_{2}|}{|\bm{\Sigma}_{1}|}-d+\text{tr}\{\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1}\}\\ &+(\bm{\mu}_{2}-\bm{\mu}_{1})^{T}\bm{\Sigma}_{2}^{-1}(\bm{\mu}_{2}-\bm{\mu}_{1})].\end{split}$ (6)
where ${\mathcal{N}}_{1}={\mathcal{N}}(\bm{\mu}_{1},\bm{\Sigma})_{1},{\mathcal{N}}_{2}={\mathcal{N}}(\bm{\mu}_{2},\bm{\Sigma}_{2})$.
@@
For this work, we set the attack target $p^{\ast}(\bm{z})$ to a zero mean Gaussian distribution $p(\bm{z})=\mathcal{N}(\mathbf{0},v\mathbf{I})$, where $v>0$ is hyper-parameter that controls the disruptive effect. Since both $p^{\ast}(\bm{z})$ and $q(\bm{z}|\bm{x})$ are diagonal multivariate Gaussian distributions, Equation 6 becomes:
$\begin{split}{\mathcal{L}}_{\text{KL}}(\bm{x})&=D_{\text{KL}}(q(\bm{z}|\bm{x})\parallel p^{\ast}(\bm{z}))\\ &=\dfrac{1}{2}\sum_{i=1}^{d}(\dfrac{\bm{\sigma}_{i}^{2}+\bm{\mu}_{i}^{2}}{v}-1+\ln(\dfrac{v}{\bm{\sigma}_{i}^{2}})).\end{split}$ (7)
Therefore, the attack formulation becomes:
$\min_{\bm{\delta}}\mathcal{L}_{\text{KL}}(\bm{x}+\bm{\delta})\quad\text{s.t.}\quad\left\lVert\bm{\delta}\right\rVert_{\infty}<\epsilon.$ (8)
Note that the attack target distribution is independent from the input image $\bm{x}$, making it a more general attack method. Successful attacks should make the encoded posterior uninformative or collapsed for a perturbed image, consequentially, the downstream image generation, i.e. Objective 2, will be disrupted accordingly as well.
Figure 3: Comparison of impact on the gradient regarding different criterion.
Note that we choose to introduce the loss $\mathcal{J}$ with $\mathit{f}$-divergence for its generality. Both forward KL and reverse KL are specific cases of the general divergence measure. For example, the reverse KL divergence is recovered when we set $\mathit{f}=\ln$. Comparing with the forward KL, the reverse KL, which is used in defining the VAE training loss and therefore also contributes to the posterior collapse issue, is found to perform better in disrupting the image generation. To see why, the reverse KL’s expectation is taken w.r.t $q(\bm{z}|\bm{x})=\mathcal{N}(\bm{\mu},\text{diag}(\bm{\sigma}^{2}))$, making the attack objective more dependent on the encoding mean and variance rather than the fixed hyperparameter $v$. Three different loss measures, mean squared error, forward KL and reverse KL, are plotted in Figure 3, where the attack variance hyperparameter $v$ is set to unit for convenience. As expected, all three losses have the same minimum: $\mu=0,\sigma^{2}=1$. However, it can be observed that the reverse KL provides a more nuanced loss compared with the symmetric MSE error. In comparison with the forward loss, reverse KL’s loss surface provides more geometric information to guide the attack, which is manifested in the steeper gradient directions across the input domain. Thus, adversary generation can be achieved by maximizing the distributional gap using gradient ascent.
To craft adversarial samples, we use projected sign gradient ascent (Madry et al. 2018) for iterative updates. Our approach is inspired by the counterintuitive finding from Xue et al. (2024), where minimizing their loss function, contrary to the intuitive maximization, yielded better results. Following this insight, we explore both gradient directions by altering the sign of our final loss function, which will be discussed in appendix. Our update process follows the form:
$\bm{x}^{t+1}=\mathcal{P}_{\ell_{\infty}\left(\bm{\delta}^{t}\right)}^{\epsilon}\left(\bm{x}^{t}+\alpha\text{sign}\nabla_{\bm{x}^{t}}\mathcal{J}\left(\bm{x}^{t}\right)\right),$ (9)
where $\mathcal{P}_{\ell_{\infty}\left(\bm{\delta}^{t}\right)}^{\epsilon}\left(\cdot\right)$ will apply the projection on $\bm{\delta}^{t}$ in the $\epsilon$-ball of $\ell_{\infty}$ norm, $\alpha$ is a hyperparameter to adjust the learning rate during the optimization.
[…]
SD14 SD15 SD14 SD15 SD14 SD15 SD14 SD15 SD14 SD15
P1 AdvDM 20.49 20.49 153.71 150.22 0.4422 0.4417 0.5010 0.5008 0.0416 0.0424
PG 18.49 18.52 151.75 147.85 0.4800 0.4840 0.5647 0.5629 0.0786 0.0792
SDS 20.60 20.56 155.55 152.76 0.4519 0.4495 0.4998 0.5007 0.0453 0.0462
Ours 18.02 18.02 139.11 135.32 0.4441 0.4453 0.6188 0.6171 0.1106 0.1110
P2 AdvDM 19.23 19.25 94.90 92.09 0.4226 0.4232 0.5017 0.5001 0.0442 0.0440
PG 17.32 17.31 88.43 91.00 0.4650 0.4660 0.5585 0.5592 0.0858 0.0865
SDS 19.32 19.31 99.80 99.03 0.4312 0.4304 0.5035 0.5029 0.0480 0.0483
Ours 16.74 16.70 70.59 69.04 0.4336 0.4329 0.6005 0.5990 0.1156 0.1167
P3 AdvDM 19.37 19.42 143.57 140.01 0.4291 0.4298 0.5142 0.5102 0.0454 0.0452
PG 17.26 17.22 123.94 127.87 0.4702 0.4749 0.5727 0.5727 0.0847 0.0853
SDS 19.56 19.59 144.63 144.64 0.4409 0.4385 0.5128 0.5102 0.0492 0.0491
Ours 16.71 16.78 156.47 153.85 0.4234 0.4228 0.6202 0.6204 0.1084 0.1079
P4 AdvDM 19.32 19.31 148.26 145.27 0.4045 0.4040 0.5347 0.5324 0.0488 0.0486
PG 17.87 17.84 150.85 151.75 0.4560 0.4589 0.5820 0.5818 0.0804 0.0819
SDS 19.49 19.48 149.37 148.82 0.4196 0.4176 0.5293 0.5275 0.0522 0.0520
Ours 17.61 17.50 116.48 115.41 0.4004 0.3942 0.6361 0.6394 0.1051 0.1067
P5 AdvDM 19.41 19.33 159.45 166.76 0.4088 0.4076 0.5239 0.5226 0.0445 0.0441
PG 17.77 17.80 110.96 112.49 0.5091 0.4988 0.5460 0.5517 0.0945 0.0924
SDS 19.48 19.39 162.63 170.77 0.4171 0.4149 0.5239 0.5236 0.0487 0.0485
Ours 16.81 16.95 134.71 129.93 0.4568 0.4524 0.6032 0.6047 0.1227 0.1209
Table 1: Comparison of our method against several baselines. Arrows ($\uparrow$/$\downarrow$) indicate whether higher or lower values represent better image quality for each image quality metric. Data is grouped vertically in sets of four for each metric and model. The best result in each group is in bold, and the second-best result is underlined.
Results
Experiments Setup
Dataset.
In our experiments, we utilized a 1000-image-subset of the ImageNet (Deng et al. 2009), as selected by Lin et al. (2020a). This choice aligns with established conventions in adversarial attack research. All images were resized to $512\times 512$, ensuring consistency across evaluations.
Baselines.
We compared our approach against several state-of-the-art methods, AdvDM (Liang et al. 2023), PhotoGuard (PG) (Salman et al. 2023), MIST (Liang and Wu 2023), and SDS (Xue et al. 2024). To ensure a fair comparison, we use the gradient ascent version of SDS, aligning with the optimization settings of other methods. We fix $t=40$ and $\epsilon=16$ for all methods, the rest left default. We implement our method with $\alpha=2$, $v=1\times 10^{-8}$, the optimization follows the goal defined in Equation 8, which the opposite one will be illustrated in Appendix.
Victim Models.
Our experiments primarily focused on popular LDMs, specifically the lightweight Stable Diffusion 1.4 (SD14) and 1.5 (SD15). These models were chosen due to their widespread accessibility and higher potential for misuse, making them relevant targets for our study on protecting against unauthorized image manipulation. To assess the transferability across different resolution and architectures, we also tested the first 100 images on Stable Diffusion 2.0 (SD20) with $768\times 768$, and Stable Diffusion XL (SDXL) with $1024\times 1024$ for all methods. All reported results of our method leverage the VAE of SD15 as the surrogate model.
Varied Prompts.
To further evaluate the robustness and versatility of our method, we conducted inference using a diverse set of prompts. These prompts include: P1: an empty prompt; P2: a caption generated by BLIP (Li et al. 2022) with the clean image; P3: “add some snow” to represent weather modifications; P4: “apply sunset lighting” to test lighting adjustments; P5: “make it like a watercolor painting” to examine style transfer capabilities. Through this varied set of prompts, we aimed to assess our method’s effectiveness across a wide range of potential image manipulation scenarios.
Image Quality Assessment (IQA).
To quantitatively assess the quality of the edited images pre- and post-attack, we employed five different IQAs. Peak Signal-to-Noise Ratio (PSNR) provides insight into the overall distortion introduced. Fréchet Inception Distance (FID) offers a measure of the perceptual quality and generation diversity (Heusel et al. 2017). Structural Similarity Index (SSIM) captures changes in local patterns of pixel intensities (Wang et al. 2004). Learned Perceptual Image Patch Similarity (LPIPS) approximately measures human perceptual similarities (Zhang et al. 2018). Aesthetic Color Distance Metric (ACDM) evaluates the overall color consistency (Guo et al. 2024b).
Against Defending Methods.
To evaluate our attack’s robustness against potential defenses, we considered two methods. First, we examined Adv-Clean [^1] https://huggingface.co/spaces/p1atdev/AdverseCleaner, a filter-based approach designed to remove adversarial perturbations in diffusion models without requiring additional training. Second, we applied $3\times 3$ Gaussian blur to simulate potential degradation of protected images during the distribution. We apply these defense on our attacked images and evaluate on SD15.
Image Editing Results with Varied Prompts
Our experimental results demonstrate the superior performance of our proposed PCA compared to existing methods. Figure 2 provides a visual comparison, while Table 1 presents quantitative results.
As observed from Figure 2, our method significantly outperforms existing approaches in terms of semantic disruption. Most other methods largely maintain an image structure similar to that of the clean image after editing. In contrast, our approach fundamentally alters the semantic information of the image, resulting in edited outputs that bear almost no resemblance to the infringer-desired images.
Even when prompts containing image structure information are provided, our method continues to demonstrate superior performance. The generated images show significant discrepancies from the expected outcomes, effectively thwarting attempts at unauthorized manipulation. Moreover, our approach introduces quite noise patterns in the generated images, substantially degrading their quality. This degradation renders the manipulated images unsuitable for their intended purposes, whether it be unauthorized use, redistribution, or further editing. Our method not only disrupts the semantic content of the image but also compromises its overall visual quality. This comprehensive degradation ensures that even if an attacker manages to partially reconstruct some elements of the original image, the result remains unusable from both a semantic and aesthetic standpoint.
The quantitative results in Table 1 further corroborate the effectiveness of our method. Notably, our Posterior Collapse Attack achieves state-of-the-art results while requiring minimal knowledge of the opponent model, less computational time, and lower GPU memory usage compared to existing methods (as illustrated in Figure 1).
These results demonstrate that, our approach effectively disrupts image semantics while maintaining efficiency and generalizability. This strategy challenges generative models’ ability to produce coherent edits, operating with minimal model-specific information in a near black-box scenario.
Transferability
As shown in Figure 4, our method demonstrates superior transferability compared to existing approaches when tested on both SD20 and SDXL models. Across all five IQA metrics, our approach consistently achieves the highest normalized scores of 1.0, indicating the best attack performance. In contrast, other methods show varying degrees of effectiveness, with PhotoGuard and MIST performing moderately well, while AdvDM and SDS exhibit limited transferability. These results validate our hypothesis regarding the limitations of current protection techniques that rely heavily on white-box access to target models. Our method, by minimizing dependence on model-specific knowledge, achieves better performance across diverse model architectures. This generalizability is crucial in addressing the challenges posed by rapidly evolving generative AI models, offering a more practical and enduring solution in this dynamic field.
[…]
Figure 5: Attack performance after Adv-Clean and JPEG degrade defense, where $\ast$ indicates the adaptive version of the method. The data in this figure under the same normalization as Figure 4 for better visualization.
Against Defense
As shown in Figure 5, our method demonstrates robust performance against the Adv-Clean defense, consistently achieving the highest normalized scores of 1.0 across all five IQA metrics. This indicates superior resilience to this type of defense compared to other methods. PhotoGuard and MIST show moderate effectiveness, while AdvDM and SDS exhibit limited resistance to Adv-Clean. Against $3\times 3$ Gaussian blur, our original method shows comparable performance, particularly in PSNR and ACDM, indicating resilience in overall image quality and color manipulation. With a simple adaptive attack technique (denoted as $\ast$), our method achieves superior performance across all metrics, consistently outperforming other approaches. This improvement demonstrates our approach’s flexibility and robustness. While the non-adaptive version has already performed well in some metrics, the adaptive version’s dominance across all metrics highlights our method’s ability to maintain effectiveness under various defenses. This adaptation allowing our method to quickly recover its original performance when under specific defense strategy. These results underscore the versatility of our attack strategy, proving its capability to overcome diverse defense mechanisms with or without minor adjustments.
Conclusion
In this paper, we proposed the Posterior Collapse Attack, a novel method for disrupting LDM-based image editing, requiring fewer parameters — merely 3.39% of the whole model, Which makes our method operates faster and consumes less VRAM than existing approaches. We leverage a novel loss function targeting the VAE encoder of LDMs, induces collapse in the VAE’s posterior distribution to significant disrupt the image semantic during LDM editing. Experimental results demonstrate that PCA significantly degrades the quality of LDM-edited images, causing a collapse in their semantic content. This approach proves highly effective in protecting images from unauthorized manipulation across various LDM architectures. Our work contributes to the ongoing efforts to secure digital assets in an era of rapidly advancing generative AI, alleviating the socio-technical challenges posed by malicious AI misuse.
Appendix
Appendix A Derivation of the Loss Function
For the Equation 6 to Equation 7 in the main paper, we provide a derivation to demonstrate how we obtained Equation 7 from Equation 6:
Given two multivariate Gaussian distribution $\mathcal{N}_{1}(\bm{\mu}_{1},\bm{\Sigma}_{1})$ and $\mathcal{N}_{2}(\bm{\mu}_{2},\bm{\Sigma}_{2})$, where $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ refer to the posterior distribution $q(\bm{z}|\bm{x})$ and our target prior distribution $p^{\ast}(\bm{z})$. We set the target attack $p^{\ast}(\bm{z})$ to a zero mean Gaussian distribution $p^{\ast}(\bm{z})=\mathcal{N}(\mathbf{0},v\mathbf{I})$. Since both $p^{\ast}(\bm{z})$ and $q(\bm{z}|\bm{x})$ are diagonal multivariate Gaussian distributions, equation can be derived as:
$D_{\text{KL}}(\mathcal{N}_{1}\parallel\mathcal{N}_{2})$ $=\frac{1}{2}\left[\ln{\frac{|\bm{\Sigma}_{2}|}{|\bm{\Sigma}_{1}|}}-d+\text{tr}\left\{\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1}\right\}+\left(\bm{\mu}_{2}-\bm{\mu}_{1}\right)^{T}\bm{\Sigma}_{2}^{-1}\left(\bm{\mu}_{2}-\bm{\mu}_{1}\right)\right]$
$=\frac{1}{2}\left(\ln{\prod_{i=1}^{d}\frac{v}{\bm{\sigma}_{i}^{2}}}-d+\sum_{i=1}^{d}\frac{\bm{\sigma}_{i}^{2}}{v}+\frac{\bm{\mu}^{T}\bm{\mu}}{v}\right)$
@@ Appendix A Derivation of the Loss Function
$D_{\text{KL}}(\mathcal{N}_{1}\parallel\mathcal{N}_{2})\simeq\mathcal{L}_{\text{KL}}(\bm{x})=\frac{1}{2}\sum_{i=1}^{d}\left(-\ln\bm{\sigma}_{i}^{2}-1+\frac{\bm{\mu}_{i}^{2}+\bm{\sigma}_{i}^{2}}{v}\right).$
Appendix B Gradient Direction
As briefly mentioned in the main paper, our investigation was motivated by the work of Xue et al. (2024), leading us to examine both the minimization (-) and maximization (+) of the loss function $\mathcal{L}_{\text{KL}}$.
To provide context for our analysis, we revisit the two primary objectives outlined in the main paper:
Objective 1: $\min_{\bm{\delta}}\mathcal{D}(f(\bm{x}+\bm{\delta}),\bm{x}+\bm{w})\,\,\,s.t.\,\|\bm{\delta}\|_{p}\leq\epsilon,$
Objective 2: $\max_{\bm{\delta}}\mathcal{D}(f(\bm{x}+\bm{\delta}),f(\bm{x}))\quad s.t.\,\|\bm{\delta}\|_{p}\leq\epsilon.$
It is crucial to note that these objectives, while seemingly different, share a common goal: to prevent the infringer-desired output. This commonality allows us to consider Objective 1 as a variant of Objective 2.
For a comprehensive evaluation of our method’s performance, we employ two distinct approaches:
  • Evaluation of Objective 1:
    We compute IQAs between clean images and edited adversarial samples, represented as $\operatorname{IQA}(\bm{x},f(\bm{x}_{adv}))$. This approach aligns with some baselines studies (Liang et al. 2023; Xue et al. 2024).
  • Evaluation of Objective 2:
    We calculate IQAs between edited clean images and edited adversarial samples, denoted as $\operatorname{IQA}(f(\bm{x}),f(\bm{x}_{adv}))$. This metric is consistent with the results presented in the main paper.
Two Variants of Posterior Collapse
Posterior collapse in VAEs is a phenomenon that can manifest in two primary forms: diffusion collapse and concentration collapse. Let $\mathcal{X}$ be the input space and $\mathcal{Z}\subseteq\mathbb{R}^{d}$ be the latent space of a VAE. The posterior distribution in a VAE is denoted by $q(\bm{z}\mid\bm{x})$, where $\bm{x}\in\mathcal{X}$ and $\bm{z}\in\mathcal{Z}$.
Here we give definitions of two kind of collapses and explain when the phenomena will occur:
  1. 1.
    Diffusion Collapse:
    where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$. In this case, the encoder produces highly uncertain latent representations for all inputs.
  2. 2.
    Concentration Collapse:
    where $\delta$ is the Dirac delta function, and $\bm{\mu}_{0}$ is typically the mean of the prior distribution. This collapse results in the encoder mapping all inputs to nearly identical latent representations.
Based on the above definitions, we propose two strategies to cause these two collapse phenomena:
@@ Appendix B Gradient Direction
    Consider the optimization problem:
@@
    When $v=1$, the KL divergence simplifies to:
    Analysis shows that maximizing this leads to:
    $|\bm{\mu}_{i}|\rightarrow\infty\quad\text{and}\quad\bm{\sigma}_{i}^{2}\rightarrow\infty\quad\text{for all }i,$ (14)
@@
    PCA+ optimization pushes the latent distribution $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$ away from the prior distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, effectively creating an out-of-distribution (OOD) scenario. This divergence introduces a significant domain shift for the downstream diffusion model, moving the data distribution into regions where the model’s knowledge is limited or inaccurate. Consequently, the diffusion model struggles to perform proper inference on the transformed data, as its fundamental assumptions about the data distribution are violated.
@@
    Consider the optimization problem:
@@
    As $v\rightarrow 0^{+}$, the dominant term becomes:

### v4

@@ IV-A Problem Definition
Objective 1: $\min_{\bm{\delta}}d(f(\bm{x}+\bm{\delta}),\bm{x})\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,s.t.\,\|\bm{\delta}\|_{p}\leq\epsilon,$
Objective 2: $\max_{\bm{\delta}}d(f(\bm{x}+\bm{\delta}),f(\bm{x}))\quad s.t.\,\|\bm{\delta}\|_{p}\leq\epsilon,$
where $f(\cdot)$ is a kind of LDM-based image editing method; $d(\cdot)$ measures the perceptual distance between two inputs; $\|\cdot\|_{p}$ applies the $\ell_{\infty}$ norm constraint to maintain visual integrity of the adversarial sample. Intuitively, Objective 1 preserves original image content by keeping the edited adversarial sample similar to $\bm{x}$, while Objective 2 disrupts unauthorized editing by maximizing deviation from the expected output $f(\bm{x})$. Both objectives prevent infringer-desired manipulations, with Objective 1 causing some distance between $f(\bm{x}+\bm{\delta})$ and $\bm{x}$ as a side effect.
Existing methods in the literature typically address either Objective 1 or Objective 2. However, the solving algorithms employed by these methods often require extensive white-box information about the target model, particularly access to the neural backbone U-Net of the LDM. This heavy reliance on model-specific details limits their transferability and applicability across different LDM architectures, requiring more computing resource.
Our method proposes a solving algorithm that can achieve both Objective 1 and Objective 2, but takes a fundamentally different approach. Instead of relying on detailed knowledge of the entire LDM pipeline, we exploit the inherent characteristics of LDM-based editing by targeting the VAE component, which is common across various LDM architectures. This strategy allows us to achieve the goal of maximizing the disparity between $f(\bm{x}_{adv})$ and $f(\bm{x})$ without requiring extensive access to model-specific information, particularly the compute-intensive and model-specific U-Net component. By concentrating on the VAE, our approach aligns more closely with real-world scenarios where full model access may not be available, providing an efficient solution to prevent infringers from exploiting LDM-based image editing outputs.
IV-B Posterior Collapse of Variational Autoencoder
A key observation driving our approach is the ubiquity of VAEs in the architecture of LDMs. VAEs serve as a foundational component across different LDM implementations, often with only minor variations between models. For instance, the architecture of VAE used in Stable Diffusion 2.0 is similar as that of Stable Diffusion 1.5. This commonality presents a strategic opportunity: by focusing on the vulnerability of VAE architecture, we can potentially affect a wide range of LDMs without requiring detailed knowledge of their specific downstream tasks.
@@ IV-B Posterior Collapse of Variational Autoencoder
By general consensus, training of VAE usually suffers an optimization issue called posterior collapsing [razavi2018preventing], that is the KL divergence term of Equation 5 dominates the overall loss such that all posteriors collapse to the uninformative prior and the reconstruction error or the likelihood term is ignored. We draw inspiration from this vulnerability and aim to propose adversarial attacks to cause the posterior collapse during inference stage.
IV-C Unified Loss Function for Two Collapse Phenomena
Building upon the posterior collapse phenomenon discussed in Section IV-B, we now present our unified approach to inducing both diffusion collapse and concentration collapse during VAE inference. Our method (shown in Fig. 2) leverages a single KL divergence-based loss function that can achieve both protection objectives through strategic parameter adjustment.
Posterior-Collapse-phenomenon in VAEs’ inference stage can manifest in two primary forms: diffusion collapse and concentration collapse. Both potential collapses could be induced in the inference stage, by maximizing and minimizing the KL divergence. Here we give definitions of two kind of collapses and explain when the phenomenon will occur. Let $\mathcal{X}$ be the input space and $\mathcal{Z}\subseteq\mathbb{R}^{d}$ be the latent space of a VAE. The posterior distribution in a VAE is denoted by $q(\bm{z}\mid\bm{x})$, where $\bm{x}\in\mathcal{X}$ and $\bm{z}\in\mathcal{Z}$,
  1. 1.
    Diffusion Collapse: (for objective 1)
    where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$. In this case, the encoder produces highly uncertain latent representations for all inputs. This can be done by maximizing the KL divergence between $\mathcal{N}(\mathbf{0},\mathbf{I})$ (i.e., setting $v=1$ in the Equation 12), which can be formulated as:
  2. 2.
    Concentration Collapse: (for objective 2)
    where $\delta$ is the Dirac delta function, and $\bm{\mu}_{0}$ is typically the mean of the prior distribution. This collapse results in the encoder mapping all inputs to nearly identical latent representations. This can be done by minimizing the KL divergence between $\mathcal{N}(\mathbf{0},\text{diag}(c))$, where $c$ is some arbitrary small number (i.e., setting $v\rightarrow 0$ in theEquation 12), and the KL divergence can be formulated as:
    $\mathcal{L}_{\mathrm{KL}}(\bm{x})=\frac{1}{2}\sum_{i=1}^{d}\left(\ln\frac{c}{\bm{\sigma}_{i}^{2}}-1+\frac{\bm{\mu}_{i}^{2}+\bm{\sigma}_{i}^{2}}{c}\right).$ (9)
@@ IV-C Unified Loss Function for Two Collapse Phenomena
Both forms can be derived from the general case, we aim to leverage the divergence measure between $q(\bm{z}|\bm{x})$ and a target distribution $p^{\ast}(\bm{z})$. The objective is to generate $\bm{x}_{adv}$ by minimising the KL divergence between $q(\bm{z}|\bm{x})$ and $p^{\ast}(\bm{z})$. The KL divergence of two multivariate Gaussian distributions is:
$\begin{split}D_{\text{KL}}(\mathcal{N}_{1}\parallel\mathcal{N}_{2})=&\frac{1}{2}\Big[\ln\frac{|\bm{\Sigma}_{2}|}{|\bm{\Sigma}_{1}|}-d+\text{tr}\{\bm{\Sigma}_{2}^{-1}\bm{\Sigma}_{1}\}\\ &+(\bm{\mu}_{2}-\bm{\mu}_{1})^{T}\bm{\Sigma}_{2}^{-1}(\bm{\mu}_{2}-\bm{\mu}_{1})\Big].\end{split}$ (10)
where $\mathcal{N}_{1}=\mathcal{N}(\bm{\mu}_{1},\bm{\Sigma}_{1}),\mathcal{N}_{2}=\mathcal{N}(\bm{\mu}_{2},\bm{\Sigma}_{2})$.
@@
Given two multivariate Gaussian distribution $\mathcal{N}_{1}(\bm{\mu}_{1},\bm{\Sigma}_{1})$ and $\mathcal{N}_{2}(\bm{\mu}_{2},\bm{\Sigma}_{2})$, where $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ refer to the posterior distribution $q(\bm{z}|\bm{x})$ and our target prior distribution $p^{\ast}(\bm{z})$. We set the target attack $p^{\ast}(\bm{z})$ to a zero mean Gaussian distribution $p^{\ast}(\bm{z})=\mathcal{N}(\mathbf{0},v\mathbf{I})$, where $v>0$ is hyper-parameter that controls the disruptive effect. Since both $p^{\ast}(\bm{z})$ and $q(\bm{z}|\bm{x})$ are diagonal multivariate Gaussian distributions, Equation 10 can be derived as:
$D_{\text{KL}}(\mathcal{N}_{1}\parallel\mathcal{N}_{2})$ $=\frac{1}{2}\left(\ln{\prod_{i=1}^{d}\frac{v}{\bm{\sigma}_{i}^{2}}}-d+\sum_{i=1}^{d}\frac{\bm{\sigma}_{i}^{2}}{v}+\frac{\bm{\mu}^{T}\bm{\mu}}{v}\right)$ (11)
@@
$D_{\text{KL}}(\mathcal{N}_{1}\parallel\mathcal{N}_{2})\simeq\mathcal{L}_{\text{KL}}(\bm{x})=\frac{1}{2}\sum_{i=1}^{d}\left(-\ln\bm{\sigma}_{i}^{2}-1+\frac{\bm{\mu}_{i}^{2}+\bm{\sigma}_{i}^{2}}{v}\right).$ (12)
Then, we choose the hyperparameter of loss function and optimization strategy in Equation 12 for both phenomena:
@@
    For objective 1, we decide to apply the diffusion collapse, which can be translate to a optimization problem:
@@
    For the KL divergence, we choose $v=1$, which align with the training process of VAE. The loss is already shown in Equation 7. Analysis shows that maximizing this leads to:
    $|\bm{\mu}_{i}|\rightarrow\infty\quad\text{and}\quad\sigma_{i}^{2}\rightarrow\infty\quad\text{for all }i,$ (14)
@@
    PCA+ optimization pushes the latent distribution $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$ away from the prior distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, effectively creating an out-of-distribution (OOD) scenario. Notably, this protection task benefits from detectability unlike evasion-focused adversarial attacks, i.e., detection cannot remove perturbations but may deter infringers, while editing disruption remains effective. This divergence introduces a significant domain shift for the downstream diffusion model, moving the data distribution into regions where the model’s knowledge is limited or inaccurate. Consequently, the diffusion model struggles to perform proper inference on the transformed data, as its fundamental assumptions about the data distribution are violated.
@@
    For objective 2, we decide to apply concentration collapse, which can be consider as a optimization problem:
@@
    For this case, we choose $v\rightarrow 0^{+}$ for the KL divergence. The loss is shown in Equation  9 and the dominant term becomes: